how do you graph the equation on a coordinate system: -24 + 3x = 7x
Answer by ewatrrr(24785)


Hi
how do you graph the following equation on a coordinate system:
-24 + 3x = 7x
-24 = 4x
-6 = x
x = -6 is a vertical line having NO slope